calculate the dynamic slope resistance when a voltage of 50V is applied across a germanium diode at 300K when saturation current is 10uA
1
Expert's answer
2020-06-15T14:45:27-0400

The current is


"I=I_s(e^{\frac{qV}{kT}}-1)."

The dynamic resistance is


"r_{ac}=\frac{kT}{qI}=\frac{kT}{qI_s(e^{\frac{qV}{kT}}-1)},\\\space\\
r_{ac}=\frac{(1.38\cdot10^{-23})300}{(1.6\cdot10^{-19})(10\cdot10^{-6})(2.72^{\frac{(1.6\cdot10^{-19})50}{(1.38\cdot10^{-23})300}}-1)}=0."

The reason for such a small resistance is the value of voltage of 50 V. Typically the value is less than 1 volt.